Find the fourth term in the expansion x^2 -2y^1/2


1
Expert's answer
2021-08-19T17:18:12-0400

"(x^2-2y)^{\frac{1}{2}}=[x^2(1-\frac{2y}{x^2})]^{\frac{1}{2}}
\\=x(1-\frac{2y}{x^2})^{\frac{1}{2}}
\\\text{Using binomial expansion, we expand the expression above}
\\=x[1-\frac{1}{2}.\frac{2y}{x^2} + \frac{\frac{1}{2}(\frac{1}{2}-1).(\frac{2y}{x^2})^2}{2!}+ \frac{\frac{1}{2}.(\frac{1}{2}-1).(\frac{1}{2}-2).(\frac{2y}{x^2})^2}{3!}]
\\=x[1-\frac{y}{x^2}-\frac{y^2}{2x^4}-\frac{y^3}{3x^6}]
\\= x-\frac{y}{x}-\frac{y^2}{2x^3}-\frac{y^3}{3x^5}+..."